what is the density of ammonia gas at 2.00 atm pressure and a temperature of 25oC


1
Expert's answer
2021-11-03T06:01:01-0400

ρ=R×TP×molarmassρ=R×TP×molar mass​

Molar mass of NH3= 17.03

= 0.082×17.03×298×2

= 832.3

= 832.23g/L